### v2

@@ 2 Related Work
Despite this growing body of work, creating real-world labeled lidar datasets remains a challenging and resource-intensive task. It demands substantial human effort, time, and financial investment. The complexity of object shapes and motions, frequent occlusions, and varying backgrounds in point cloud data—particularly from the roadside perspective—make the annotation process labor-intensive and highly dependent on skilled expertise.
@@ 3.1 Generation Approach
Dynamic Content and Annotation: The dynamic elements in our datasets, such as types of road-users (cars, trucks, cyclists, etc.), are generated stochastically, and therefore, do not reproduce the real trajectories of the real-world actors. However, their behavior is based on the traffic rules and physics based interactions modeled in our custom maps for CARLA. The simulations include randomized but realistic traffic patterns introducing contrast in object positions, motions, and occlusion scenarios. Each frame is annotated with 3D bounding-boxes, object IDs, and semantic tags supporting tasks such as object detection, tracking, and scene understanding.
@@ 3.2 Dataset Description
Each synthetic dataset contains 10K frames of intersection focused activity, covering roughly the same area as the target datasets. The datasets include point-cloud data along with rich annotations. An overview of the datasets is presented in Table 1.
@@ 6 Conclusion & Future Work
This paper introduced three high-fidelity synthetic roadside lidar datasets, UT-LUMPI, UT-V2X-Real, and UT-TUMTraf-I, constructed using a digital-twin-based pipeline to replicate the physical layout and sensor properties of real-world deployments. Unlike generic synthetic datasets, our approach achieves close structural, statistical, and sensory alignment with real data. Through distributional analysis and object detection case studies, we demonstrated that models trained solely on synthetic data generalize well to real-world benchmarks, in some cases outperforming models trained on human-annotated datasets. These results highlight the utility of UrbanTwin datasets as both augmentation and standalone training sources. By bridging the sim-to-real gap in roadside lidar perception, UrbanTwin offers a scalable, cost-effective resource for both research and deployment in intelligent transportation systems. The modular simulation framework supports extensibility to new classes, environments, and annotation types. Future work will focus on integrating pedestrian and other Vulnerable Road User (VRU) classes, whose inclusion is critical for safety-related ITS research [47]. To overcome current simulation limitations, we plan to integrate realistic human models using technologies like Unreal Engine’s MetaHuman framework [48] and source naturalistic motion from large-scale motion capture libraries (e.g. AMASS [49]). We will also explore cross-site generalization, expanding validation across tracking, segmentation, and sensor fusion tasks, and continue creating synthetic replicas for additional public lidar datasets as part of the UCF Digital Twin initiative.

### v3

@@ 2 Related Work
Despite this growing body of work, creating real-world labeled lidar datasets remains a challenging and resource-intensive task. It demands substantial human effort, time, and financial investment. The complexity of object shapes and motions, frequent occlusions, and varying backgrounds in point cloud data, particularly from the roadside perspective, make the annotation process labor-intensive and highly dependent on skilled expertise.
@@ 3.1 Generation Approach
Dynamic Content and Annotation: The dynamic elements in our datasets are generated stochastically. While they do not reproduce the exact historical trajectories (trajectory replay) of the real-world actors (e.g., cars, trucks, cyclists, etc.), their behavior is constrained by the real-world lane topology, traffic signal phases, and physics-based interactions modeled in our custom CARLA maps. This stochastic approach serves as a form of data augmentation, generating unique interaction scenarios that statistically conform to the real site while providing scene diversity. Each generated point cloud frame is annotated with 3D bounding-boxes, object IDs, and semantic tags supporting tasks such as object detection, tracking, and scene understanding.
@@ 3.2 Dataset Description
Each synthetic dataset contains 10K frames of intersection focused activity, covering roughly the same area as the target datasets. The datasets include point cloud data along with rich annotations. An overview of the datasets is presented in Table 1.
@@ 6 Conclusion & Future Work
This paper introduced three high-fidelity synthetic roadside lidar datasets, UT-LUMPI, UT-V2X-Real, and UT-TUMTraf-I, constructed using a digital-twin-based pipeline to replicate the physical layout and sensor properties of real-world deployments. Unlike generic synthetic datasets, our approach achieves close structural, statistical, and sensory alignment with real data. Through distributional analysis and object detection case studies, we demonstrated that models trained solely on synthetic data generalize well to real-world benchmarks, in some cases outperforming models trained on human-annotated datasets. These results highlight the utility of UrbanTwin datasets as both augmentation and standalone training sources. By bridging the sim-to-real gap in roadside lidar perception, UrbanTwin offers a scalable, cost-effective resource for both research and deployment in intelligent transportation systems.
While UrbanTwin achieves strong geometric alignment, we acknowledge limitations to be addressed in future. First, strictly stochastic traffic generation, while excellent for increasing sample diversity, may not fully capture the complex, naturalistic interactions (e.g., yielding behavior or aggressive maneuvering) seen in real conflict zones. Future work will explore integrating data-driven behavioral models to enhance the realism of agent dynamics while maintaining the generative flexibility of the simulation. Second, pedestrians, whose inclusion is critical for safety-related ITS research [46], were omitted from this release to ensure high fidelity, as standard simulator assets often lack the geometric realism required for Lidar training. We are currently working to integrate realistic human models using Unreal Engine’s MetaHuman framework [11] and naturalistic motion from the AMASS library [23] to close this gap. Finally, this release focuses on clear weather conditions to establish a baseline for structural fidelity. While the simulation framework supports environmental changes, accurate modeling of Lidar physics in adverse weather (rain, fog) remains a challenge. Future updates will leverage this Digital Twin framework to introduce validated weather conditions, as well as variations in sensor placement and traffic density, to further their generalizability and applicability. We will also expand validation across tracking, segmentation, and sensor fusion tasks, and continue creating synthetic replicas for additional public lidar datasets as part of the UCF Digital Twin initiative.